# LLM-Enhanced Multi-Agent Reinforcement Learning with Expert Workflow for Real-Time P2P Energy Trading

### v2

@@ V CONCLUSION
This paper addresses the collaborative decision-making challenges among multiple prosumers in local real-time P2P electricity markets by proposing a framework that integrates LLM expert guidance with multi-agent imitation learning. This approach effectively overcomes limitations inherent in traditional optimization methods–particularly their inability to achieve real-time decision-making–as well as limitations of MARL without LLM guidance, particularly high manual labor costs associated with human expert involvement. The framework innovatively employs LLMs as expert to generate personalized strategies for guiding MARL training, combined with Wasserstein metric and an enhanced Critic network, achieving deep integration of expert knowledge and agent learning. This significantly reduces manual costs while enhancing policy optimization performance.

### v3

@@ V CONCLUSION
This paper addresses the collaborative decision-making challenges among multiple prosumers in local real-time P2P electricity markets by proposing a framework that integrates LLM expert guidance with MARL. This approach effectively overcomes limitations inherent in traditional optimization methods–particularly their inability to achieve real-time decision-making–as well as limitations of MARL without LLM guidance, particularly high manual labor costs associated with human expert involvement. The framework innovatively employs LLMs as experts to generate personalized strategies for guiding MARL training, combined with the Wasserstein metric and an enhanced Critic network, achieving deep integration of expert knowledge and agent learning. This significantly reduces manual costs while enhancing policy optimization performance.
@@
Computationally, the framework’s overhead is concentrated in the training stage, while real-time operation remains highly efficient using lightweight networks. The highest average cost of the LLM expert workflow is less than 6.27 CNY. Retraining is infrequent, triggered only by structural changes like topology modification or market adjustments, rather than routine fluctuations. Crucially, the training duration is negligible compared to the long intervals of these structural evolutions. Furthermore, the system supports efficient continual learning and knowledge transfer; the shared critic preserves learned network constraints, and the modular code repository enables direct reuse, significantly reducing redundancy and ensuring scalability.